### v2

# MT2ST: Adaptive Multi-Task to Single-Task Learning

Abstract
The conventional training approaches often face challenges in balancing the breadth of multi-task learning (MTL) with the depth of single-task learning (STL). To address this issue, we introduce the Multi-Task to Single-Task (MT2ST) framework, a groundbreaking approach that can combine the generalizability of MTL with the precision of STL. Our work include two strategies: ’Diminish’ and ’Switch’. ’Diminish’ Strategy will gradually reduce the influence of auxiliary tasks, while the ’Switch’ strategy involves a shift from multi-tasking to single-tasking at a specific timepoint at the training process.
In this paper, we propose the Multi-Task to Single-Task (MT2ST) framework, a novel approach that significantly enhances the efficiency and accuracy of word embedding training while concurrently addressing prevalent issues such as overfitting. Our empirical studies demonstrate that MT2ST can reduce training time by 67% when contrasted with single-task learning approaches, and by 13% compared to traditional multi-task learning methods. These findings underscore MT2ST’s potential to be a powerful tools for word embedding training acceleration. The code implementation is can be found at: https://github.com/NoakLiu/MT2ST-Word-Embeddings-Acceleration.
Index Terms:
Efficient ML, ML Acceleration
I Introduction
Recent developments in natural language processing have highlighted that word embeddings are foundational to many NLP tasks,its performance and training efficiency directly impacting the effectiveness of linguistic tasks.
Traditional word embedding training methods predominantly rely on either MTL, which offers a wide learning scope at the expense of task-specific detail, or STL, which excels in focus on specific task but lacks the comprehensiveness of MTL. [ZWX+24]. Our research introduces a novel paradigm, the MT2ST framework, which combines comprehensive scope of MTL and task-sepecific precision of STL together. [CPKC09]
The MT2ST framework incorporates two specific strategies for training word embeddings: ’Diminish’ and ’Switch’. The Diminish strategy innovatively reduces the impact of auxiliary tasks over time, ensuring a gradual shift to STL. In contrast, the ’Switch’ strategy splits training into two stages: starting with MTL and then making a direct transition to STL. MT2ST framework has been proven effectiveness in our emprical study. [BOHG19]
Our contributions can be summarized as follows:
  • Bridge the gap between MTL and STL with the unique MT2ST framework.
  • Introduce two pioneering strategies within this framework. ’Diminish’ and ’Switch’.
  • Empirically demonstrate MT2ST’s superiority in training efficiency and model performance. [SK18]
II Motivation
II-A Challenges in Single-Task Word Embedding Training
Word embeddings map words to continuous vector spaces, reflecting their semantic meanings. It has wide applications in text classification, sentiment analysis, and machine translation. [AX23] However, its training process is hindered by several constraints, which are as follows:
Data Volume Requirement: The need for large-scale textual datasets to train word embeddings results in significant time and computational resource consumption. Processing these large datasets is inherently slow and demands extensive computational power. [EWVD19]
Computational Complexity: Training involves complex mathematical computations. These algorithms, while effective, require high computational resources, leading to increased training duration and processing power needs. Some recent work such as [LWJV24], [LJP24] has proven the importance of efficient computing in reducing computational resources through algorithm and system co-design optimization.
Optimization Difficulties: Optimizing word embedding models for accurate semantic representation is a meticulous process. It often necessitates iterative tuning and adjustments, which are resource-intensive and time-consuming.
II-B Optimizing Training Efficiency in NLP Models with Multi-Task Learning
The training efficiency of word embeddings is hindered by the mismatch between large training datasets and narrow objective function metrics. This often leads to underutilization of datasets and focuses on limited semantic or syntactic aspects, causing inefficiencies. Altough work like attention based multi-task learning [LJD19] can adjust weights for different task overtim, models still struggle with polysemous words due to objective functions missing their subtle meanings, leading to repetitive training and reduced efficiency.
Multi-task learning (MTL) emerges as an effective solution for these inefficiencies. [ZYY+23] Recent research indicates that training on multiple related tasks can yield computational benefits. MTL has wide application in industry, such as building end-to-end dialogue modelling. [SSM+22] MTL enables models to share core representations across tasks, enhancing word embeddings with a richer semantic understanding, thus improving efficiency.
II-C Potential Inefficiency of Multi-Task Learning in Word Embedding Training
MTL brings notable benefits but can also introduce inefficiencies, especially in word embedding training. In MTL, the shared embedding layer across tasks may receive divergent gradient directions from different tasks, which may conflict.
This issue arises as tasks emphasize distinct aspects of word meanings. For example, sentiment analysis prioritizes emotional connotations, while part-of-speech tagging focuses on grammatical features. Consequently, the embedding layer faces mixed directions of gradient descent in tuning embeddings, which can potentially result in less effective task-specific embeddings. [LLJ+21]
Additionally, computational demands rise in MTL. Each task’s will update the parameters the embedding layer, which leads to time inefficiency and high resource usage. This issue intensifies with large vocabularies and high-dimensional embeddings common in NLP.
III Model
III-A MT2ST Framework
In this research, we propose ”MT2ST” (Multi-Task to Single-Task) model to improve the efficiency of word embedding training, specifically addressing the challenges in multi-task learning (MTL) environments.
The process begins with a multi-task learning phase, where a unified model, featuring a shared embedding layer, is simultaneously trained across various NLP tasks. This phase utilizes the advantages of MTL, enabling the model to absorb a wide array of linguistic features and nuances. The shared embedding layer benefits from varied linguistic inputs, fostering a more comprehensive word representation.
Once the multi-task phase is established, MT2ST shifts focus to single-task learning. Here, the pre-trained shared embeddings are fine-tuned in a task-specific manner. This targeted fine-tuning aligns the embeddings more closely with the distinct demands and subtleties of each task, ensuring that while they maintain the general linguistic knowledge from the MTL phase, they are also tailored for superior performance in specific tasks.
MT2ST employs advanced techniques for a smooth and effective transition from multi-task to single-task learning, which provides a fine-tuning and pretraining method for effcient NLP[TLJ+23]. These include adaptive learning rate mechanisms and selective freezing of certain embedding vector dimensions based on their relevance to the specific task. These strategies enable a balance between retaining the broad linguistic understanding gained in the MTL phase and focusing on task-specific details during the single-task phase.
III-B Model Construction
Instead of mono-task fine-tuning used in Multi-Task Supervised Pretraining [MST+20], our MT2ST model develop an adaptive method to accelerate word embedding training within a multi-task to single-task learning framework.
III-B1 Input Layer
  • Processes raw text input.
  • Converts text into a numerical format suitable for NLP tasks using tokenization and encoding techniques.
III-B2 Embedding Layer
  • Learns dense vector representations of words, capturing semantic and syntactic information.
  • Trained across multiple tasks in the initial phase to obtain generalized word embeddings.
III-B3 Shared Layers
  • Extract and learn features common across different tasks.
  • Play a crucial role in the multi-task learning phase.
III-B4 Dynamic Weighting Mechanism (Dimish Strategy Only)
  • Integral to the Diminish strategy, this mechanism adjusts the influence of each task over time.
  • Employs time-dependent weighting functions that reduce the impact of auxiliary task losses as the training progresses.
  • Ensures that the transition from multi-task learning to single-task learning is smooth and performance-driven.
III-B5 Task-Specific Layers (Switch Strategy Only)
  • Dedicated layers for each task, activated during the single-task learning phase in the Switch strategy.
  • Fine-tuned to align the model’s output closely with the requirements of the specific task.
III-B6 Feedback Loop
  • Continuously evaluates the model’s performance, providing insights for adjusting the dynamic weighting and learning parameters.
  • Facilitates a data-driven approach to optimize the transition phase in the training process.
III-C Model Overview
Figure 1: MT2ST Word Embedding Training Framework Overview22footnotemark: 2
III-D MT2ST: Diminish Strategy
The Diminish strategy integrates a time-dependent weighting mechanism to progressively reduce the influence of auxiliary tasks, thereby increasingly highlight the primary task i the training process.
Figure 2: MT2ST-Dimish: Dimish Strategy for Auxiliary Tasks in Training to Single Task Learning
Algorithmic Implementation: Given an input sequence $X$, embeddings are obtained via an encoder network (such as BiLSTM):
$Emb=Encoder(X;\theta_{Encoder})$ (1)
where $\theta_{Encoder}$ represents the parameters of the encoder. The gradients for both the primary task $\nabla\mathcal{L}_{primary}$ and the auxiliary tasks $\nabla\mathcal{L}_{aux_{k}}$ are computed. The update of the model parameters at each epoch $t$ is calculated by a diminishing weight for each auxiliary task, defined as:
$\gamma_{k}(t)=\gamma_{k,0}\cdot e^{-\eta_{k}t^{\nu_{k}}}$ (2)
$\gamma_{k,0}$ is the initial weight, $\eta_{k}$ is the decay rate, and $\nu_{k}$ is the curvature parameter of the decay function for the $k$-th auxiliary task. The parameter update rule is as follow:
$\theta^{(t+1)}=\theta^{(t)}-\eta\left(\nabla\mathcal{L}_{primary}+\sum_{k=1}^{K}\gamma_{k}(t)\nabla\mathcal{L}_{aux_{k}}\right)$ (3)
The MT2ST-Diminish gradient descent iterates until the primary task loss converges, which indicates the model’s gradually focus from multi-task to primary-task.
Figure 3: MT2ST to realize quicker convergence [An example of MT2ST Diminish]
input : Initial parameters $\theta^{(0)}$, learning rates $\{\eta_{k}\}_{k=1}^{K}$, decay curvature parameters $\{\nu_{k}\}_{k=1}^{K}$, sequence input $X$
output : Optimized parameters $\theta^{*}$
initialize : $\theta\leftarrow\theta^{(0)}$; $\gamma_{k}\leftarrow\gamma_{k,0}$ for $k\in\{1,\ldots,K\}$
for $t\leftarrow 1$ to $T$ do
       $Emb\leftarrow$ Encoder($X,\theta$);
      $\nabla\mathcal{L}_{primary}\leftarrow$ BackwardPass($Emb,\theta$, primary);
      for $k\leftarrow 1$ to $K$ do
             $\nabla\mathcal{L}_{aux_{k}}\leftarrow$ BackwardPass($Emb,\theta$, $aux_{k}$);
            $\gamma_{k}(t)\leftarrow\gamma_{k,0}\cdot e^{-\eta_{k}t^{\nu_{k}}}$;
       end for
      $\Delta\theta\leftarrow\nabla\mathcal{L}_{primary}$;
      for $k\leftarrow 1$ to $K$ do
             $\Delta\theta\leftarrow\Delta\theta+\gamma_{k}(t)\cdot\nabla\mathcal{L}_{aux_{k}}$;
       end for
      $\theta\leftarrow\theta-\eta\cdot\Delta\theta$;
      if CheckConvergence($\nabla\mathcal{L}_{primary}$) then
             return $\theta$;
       end if
end for
Algorithm 1 Diminish Strategy for MT2ST
III-E MT2ST: Switch Strategy
The Switch strategy features a clear transition in its training approach. Initially, the model simultaneously engages with both primary and auxiliary tasks. However, upon reaching a predetermined epoch, denoted as $T_{\text{switch}}$, the model shifts its focus solely to the primary task.
Figure 4: MT2ST-Switch: Optimal Solution Area Searching and Task Switching
Algorithmic Implementation: The key difference between Switch and Diminish strategy is the parameter update methodology. For epochs before $T_{switch}$, the parameter update is a combination of the gradients from all tasks:
$\theta^{(t+1)}=\theta^{(t)}-\eta\left(\nabla\mathcal{L}_{primary}+\sum_{k=1}^{K}\nabla\mathcal{L}_{aux_{k}}\right),\text{ for }t<T_{switch}$ (4)
Post $T_{switch}$, the update focuses solely on the primary task:
$\theta^{(t+1)}=\theta^{(t)}-\eta\nabla\mathcal{L}_{primary},\text{ for }t\geq T_{switch}$ (5)
This strategy leverages the initial shared learning from multiple tasks and then refines the model’s capabilities to focus more intently on the primary task by switch.
input : Initial parameters $\theta^{(0)}$, switch epoch $T_{switch}$, sequence input $X$
output : Optimized parameters $\theta^{*}$
initialize : $\theta\leftarrow\theta^{(0)}$
for $t\leftarrow 1$ to $T$ do
       $Emb\leftarrow$ Encoder($X,\theta$);
      if $t<T_{switch}$ then
             $\nabla\mathcal{L}_{primary}\leftarrow$ BackwardPass($Emb,\theta$, primary);
             for $k\leftarrow 1$ to $K$ do
                   $\nabla\mathcal{L}_{aux_{k}}\leftarrow$ BackwardPass($Emb,\theta$, $aux_{k}$);
             end for
            $\Delta\theta\leftarrow\nabla\mathcal{L}_{primary}+\sum_{k=1}^{K}\nabla\mathcal{L}_{aux_{k}}$;
      else
             $\nabla\mathcal{L}_{primary}\leftarrow$ BackwardPass($Emb,\theta$, primary);
            $\Delta\theta\leftarrow\nabla\mathcal{L}_{primary}$;
       end if
      $\theta\leftarrow\theta-\eta\cdot\Delta\theta$;
      if CheckConvergence($\nabla\mathcal{L}_{primary}$) then
             return $\theta$;
       end if
end for
Algorithm 2 Switch Strategy for MT2ST
IV Result Analysis
To evaluate the performance of MT2ST, various experimental setups were implemented (experimental settings are available in the appendix).
In the first experimental group, the approach was to directly fine-tune the pre-pretrained embeddings on a dataset dedicated to a single task. Observations from this experiment indicated a steady decrease in Loss, but there were fluctuations in accuracy, indicating overfitting in the model.
Figure 5: Accuracy and Loss change of Original Mode
In the second experimental series, a multi-task learning approach was applied. This approach led to a consistent improvement in both Loss and accuracy metrics. The model achieved a notable accuracy rate of 68% on the test set.
Figure 6: Loss change (Multi-Task Learning)
Figure 7: Accuracy change (Multi-Task Learning)
In the third set of experiments, the MT2ST-diminish strategy was employed. Unlike previous tests, the Loss did not decrease uniformly, which was attributed to the reduction of auxiliary task Loss during the training process. However, the model’s accuracy consistently improved, ultimately achieving a higher test set accuracy of 71% compared to the standard multi-task learning approach. This result highlights the effectiveness of the MT2ST-diminish strategy in facilitating a smooth transition from multi-task to task-specific learning, thereby enhancing the model’s overall performance.
Figure 8: Loss change in Mode (MT2ST: Dimish)
Figure 9: Accuracy change (MT2ST: Dimish)
The fourth experimental group implemented fine-tuning on the dataset using predefined embeddings through the MT2ST-switch strategy. This approach consisted of two phases: initial multi-task learning followed by task-specific learning, with an abrupt transition between the two. As indicated by the red line in the graph, a step change in Loss occurred after the switch, yet there was a steady improvement in both phases. This method achieved a high score of 77% on the test set, demonstrating its effectiveness.”
Figure 10: Loss change (MT2ST: Switch)
Figure 11: Accuracy change (MT2ST: Swtich)
V Discussion
Our empirical experiments reveal that the MT2ST model notably outperforms conventional embedding training methods in terms of training efficiency and overall quality. The newly developed embeddings exhibit superior performance and robustness across a range of downstream NLP tasks, including sentiment analysis, part-of-speech tagging, and named entity recognition. These results underscore the effectiveness of MT2ST in reducing training times and lessening computational resource requirements. The multi-task to single task learning also has applications in some recent work, such as [WCL24] [LWC+24] [XQW+24] [DLL24]. This efficiency makes MT2ST a valuable solution for fast training and high-quality word representations.