### v3

# MT2ST: Adaptive Multi-Task to Single-Task Learning

Abstract
Efficient machine learning (ML) has become increasingly important as models grow larger and data volumes expand. In this work, we address the trade-off between generalization in multi-task learning (MTL) and precision in single-task learning (STL) by introducing the Multi-Task to Single-Task (MT2ST) framework. MT2ST is designed to enhance training efficiency and accuracy in word embedding tasks, showcasing its value as a practical application of efficient ML.
Our framework employs two strategies: *Diminish*, which gradually reduces the influence of auxiliary tasks, and *Switch*, which transitions training from MTL to STL at a specific point. Empirical results show that MT2ST reduces training time by 67% compared to STL and by 13% compared to traditional MTL, while maintaining high accuracy. These findings highlight MT2ST as an efficient ML solution tailored for optimizing word embedding training. Code is available at https://github.com/NoakLiu/MT2ST.
Index Terms:
Efficient ML, Model Compression, ML Acceleration
I Introduction
Efficient machine learning (ML) has become increasingly crucial due to the growing size of models and datasets across various domains such as natural language processing (NLP), computer vision, and speech recognition. The need for efficient ML has driven innovations in areas such as embedding generation for semantic representation, multi-task learning (MTL) for generalization across related tasks, and optimization techniques that reduce computational demands. However, the training of embeddings faces several challenges, including computational inefficiencies, optimization difficulties, and scalability issues, as highlighted by Mikolov et al. (2013) [MCCD13] and more recent surveys on NLP efficiency [Z+23].
Embedding training typically utilizes either single-task learning (STL) or multi-task learning (MTL). STL is known for providing precise task-specific learning but lacks the broader generalization capabilities of MTL, as discussed by Collobert et al. (2011) [CW11]. On the other hand, MTL enables models to share representations across tasks, enhancing generalization. However, it introduces challenges such as gradient conflicts and inefficiencies in resource usage, particularly when dealing with tasks that require distinct representations [SK18, Z+23]. This makes balancing the breadth of MTL with the precision of STL an ongoing challenge in machine learning.
In this paper, we introduce the Multi-Task to Single-Task (MT2ST) framework, which addresses this challenge by combining the strengths of both MTL and STL. MT2ST employs two novel strategies: Diminish and Switch. The Diminish strategy gradually reduces the influence of auxiliary tasks over time, ensuring a smooth transition to STL, while the Switch strategy shifts the model from multi-tasking to single-tasking at a predefined point. Our framework significantly improves training efficiency and accuracy, as demonstrated in our empirical studies. Specifically, we show that MT2ST reduces training time by 67% compared to STL and by 13% compared to traditional MTL, while maintaining high accuracy and model performance.
The contributions of our work can be summarized as follows:
  • Bridging the gap between MTL and STL with the innovative MT2ST framework.
  • Introducing the Diminish and Switch strategies for efficient learning.
  • Demonstrating MT2ST’s superior training efficiency and performance in word embedding tasks.
II Motivation
II-A Challenges in Single-Task Embedding Training
Embeddings are widely used for representing objects, such as words, images, or even users, in continuous vector spaces that capture their semantic relationships. Despite their broad applicability across domains, embedding generation faces several challenges:
Data Volume: Large-scale datasets are often required for training high-quality embeddings, resulting in significant time and computational resource consumption. Processing these large datasets is inherently slow and computationally expensive [EWVD19, LJP24].
Computational Complexity: The generation of embeddings involves complex mathematical computations, demanding high computational power and time. Efficient computational methods are essential to optimize these processes. Recent work has focused on algorithmic optimizations and system co-design to reduce computational demands, such as in [LWJV24, LJP24].
Optimization Difficulties: Training embeddings for accurate semantic or contextual representation is a complex process, requiring iterative optimization. This optimization is resource-intensive and time-consuming, particularly when dealing with large datasets or high-dimensional embeddings. Research such as [G+21, S+21] highlights the challenges in optimizing embeddings while maintaining their quality across different tasks.
II-B Optimizing Training Efficiency with Multi-Task Learning
Embedding generation can be made more efficient by using multi-task learning (MTL), where models are trained on related tasks simultaneously. This allows for the sharing of representations across tasks, enhancing the quality of the embeddings by leveraging additional, related information. MTL has been widely applied in various domains, such as NLP, computer vision, and speech recognition. For example, in computer vision, multi-task networks improve object detection and segmentation by learning shared features across related tasks [M+16, Z+20]. In NLP, MTL improves contextual understanding of embeddings [ZYY+23, SSM+22].
Recent work has shown that MTL can significantly improve efficiency, not only in NLP but also in other domains like computer vision or speech recognition, where it is applied to tasks such as object detection or speaker identification. MTL enables models to generalize across multiple tasks, leading to more robust representations [G+21].
II-C Inefficiencies in Multi-Task Learning for Embedding Generation
Despite its benefits, MTL can also introduce inefficiencies, especially in the context of embedding generation. One issue arises when different tasks impose conflicting gradients on the shared embedding representation. For instance, while one task might prioritize fine-grained details, another might emphasize broader patterns, leading to conflicting gradient updates that can hinder the optimization process. This conflict is further discussed in recent works such as [LLJ+21].
Furthermore, MTL can be computationally expensive, as each task requires updates to the shared embedding layer, resulting in increased resource consumption and longer training times. This is particularly problematic when working with large-scale datasets or high-dimensional embeddings. Work by [liu2020multi] and [SSM+22] has examined the computational costs and the challenges that arise with increasing task complexity.
III Model
III-A MT2ST Framework
We introduce the MT2ST (Multi-Task to Single-Task) framework to optimize embedding generation training. It combines multi-task learning (MTL) and single-task learning (STL) to achieve efficient training while overcoming common challenges in multi-task environments.
The process starts with MTL, where a unified model with a shared embedding layer is trained across multiple tasks. This allows the model to capture diverse linguistic features and semantic knowledge. The shared embedding layer benefits from varied inputs, providing a more generalized word representation [LJD19].
After the MTL phase, MT2ST transitions to STL, fine-tuning the pre-trained embeddings for specific tasks. This phase refines the embeddings to match the unique requirements of each task, improving performance while retaining the knowledge gained from the MTL phase. Techniques like adaptive learning rates and selective freezing of embedding dimensions ensure a smooth transition and maintain the balance between generalization and specialization [TLJ+23].
III-B Model Construction
Instead of mono-task fine-tuning used in Multi-Task Supervised Pretraining [MST+20], our MT2ST model develop an adaptive method to accelerate embedding generation through a multi-task to single-task learning framework.
III-B1 Input Layer
  • Processes raw text input.
  • Converts text into a numerical format suitable for NLP tasks using tokenization and encoding techniques.
III-B2 Embedding Layer
  • Learns dense vector representations of embedding, capturing semantic and syntactic information.
  • Trained across multiple tasks in the initial phase to obtain generalized word embeddings.
III-B3 Shared Layers
  • Extract and learn features common across different tasks.
  • Play a crucial role in the multi-task learning phase.
III-B4 Dynamic Weighting Mechanism (Dimish Strategy Only)
  • Integral to the Diminish strategy, this mechanism adjusts the influence of each task over time.
  • Employs time-dependent weighting functions that reduce the impact of auxiliary task losses as the training progresses.
  • Ensures that the transition from multi-task learning to single-task learning is smooth and performance-driven.
III-B5 Task-Specific Layers (Switch Strategy Only)
  • Dedicated layers for each task, activated during the single-task learning phase in the Switch strategy.
  • Fine-tuned to align the model’s output closely with the requirements of the specific task.
III-B6 Feedback Loop
  • Continuously evaluates the model’s performance, providing insights for adjusting the dynamic weighting and learning parameters.
  • Facilitates a data-driven approach to optimize the transition phase in the training process.
III-C Model Overview
Figure 1: MT2ST Embedding Training Framework Overview
III-D MT2ST: Diminish Strategy
The Diminish strategy integrates a time-dependent weighting mechanism to progressively reduce the influence of auxiliary tasks, thereby increasingly highlight the primary task i the training process.
Algorithmic Implementation: Given an input sequence $X$, embeddings are obtained via an encoder network (such as BiLSTM):
$Emb=Encoder(X;\theta_{Encoder})$ (1)
where $\theta_{Encoder}$ represents the parameters of the encoder. The gradients for both the primary task $\nabla\mathcal{L}_{primary}$ and the auxiliary tasks $\nabla\mathcal{L}_{aux_{k}}$ are computed. The update of the model parameters at each epoch $t$ is calculated by a diminishing weight for each auxiliary task, defined as:
$\gamma_{k}(t)=\gamma_{k,0}\cdot e^{-\eta_{k}t^{\nu_{k}}}$ (2)
$\gamma_{k,0}$ is the initial weight, $\eta_{k}$ is the decay rate, and $\nu_{k}$ is the curvature parameter of the decay function for the $k$-th auxiliary task. The parameter update rule is as follow:
$\theta^{(t+1)}=\theta^{(t)}-\eta\left(\nabla\mathcal{L}_{primary}+\sum_{k=1}^{K}\gamma_{k}(t)\nabla\mathcal{L}_{aux_{k}}\right)$ (3)
The MT2ST-Diminish gradient descent iterates until the primary task loss converges, which indicates the model’s gradually focus from multi-task to primary-task.
input : Initial parameters $\theta^{(0)}$, learning rates $\{\eta_{k}\}_{k=1}^{K}$, decay curvature parameters $\{\nu_{k}\}_{k=1}^{K}$, sequence input $X$
output : Optimized parameters $\theta^{*}$
initialize : $\theta\leftarrow\theta^{(0)}$; $\gamma_{k}\leftarrow\gamma_{k,0}$ for $k\in\{1,\ldots,K\}$
for $t\leftarrow 1$ to $T$ do
       $Emb\leftarrow$ Encoder($X,\theta$);
      $\nabla\mathcal{L}_{primary}\leftarrow$ BackwardPass($Emb,\theta$, primary);
      for $k\leftarrow 1$ to $K$ do
             $\nabla\mathcal{L}_{aux_{k}}\leftarrow$ BackwardPass($Emb,\theta$, $aux_{k}$);
            $\gamma_{k}(t)\leftarrow\gamma_{k,0}\cdot e^{-\eta_{k}t^{\nu_{k}}}$;
       end for
      $\Delta\theta\leftarrow\nabla\mathcal{L}_{primary}$;
      for $k\leftarrow 1$ to $K$ do
             $\Delta\theta\leftarrow\Delta\theta+\gamma_{k}(t)\cdot\nabla\mathcal{L}_{aux_{k}}$;
       end for
      $\theta\leftarrow\theta-\eta\cdot\Delta\theta$;
      if CheckConvergence($\nabla\mathcal{L}_{primary}$) then
             return $\theta$;
       end if
end for
Algorithm 1 Diminish Strategy for MT2ST
III-E MT2ST: Switch Strategy
The Switch strategy features a clear transition in its training approach. Initially, the model simultaneously engages with both primary and auxiliary tasks. However, upon reaching a predetermined epoch, denoted as $T_{\text{switch}}$, the model shifts its focus solely to the primary task.
input : Initial parameters $\theta^{(0)}$, switch epoch $T_{switch}$, sequence input $X$
output : Optimized parameters $\theta^{*}$
initialize : $\theta\leftarrow\theta^{(0)}$
for $t\leftarrow 1$ to $T$ do
       $Emb\leftarrow$ Encoder($X,\theta$);
      if $t<T_{switch}$ then
             $\nabla\mathcal{L}_{primary}\leftarrow$ BackwardPass($Emb,\theta$, primary);
             for $k\leftarrow 1$ to $K$ do
                   $\nabla\mathcal{L}_{aux_{k}}\leftarrow$ BackwardPass($Emb,\theta$, $aux_{k}$);
             end for
            $\Delta\theta\leftarrow\nabla\mathcal{L}_{primary}+\sum_{k=1}^{K}\nabla\mathcal{L}_{aux_{k}}$;
      else
             $\nabla\mathcal{L}_{primary}\leftarrow$ BackwardPass($Emb,\theta$, primary);
            $\Delta\theta\leftarrow\nabla\mathcal{L}_{primary}$;
       end if
      $\theta\leftarrow\theta-\eta\cdot\Delta\theta$;
      if CheckConvergence($\nabla\mathcal{L}_{primary}$) then
             return $\theta$;
       end if
end for
Algorithm 2 Switch Strategy for MT2ST
Algorithmic Implementation: The key difference between Switch and Diminish strategy is the parameter update methodology. For epochs before $T_{switch}$, the parameter update is a combination of the gradients from all tasks:
$\theta^{(t+1)}=\theta^{(t)}-\eta\left(\nabla\mathcal{L}_{primary}+\sum_{k=1}^{K}\nabla\mathcal{L}_{aux_{k}}\right),\text{ for }t<T_{switch}$ (4)
Post $T_{switch}$, the update focuses solely on the primary task:
$\theta^{(t+1)}=\theta^{(t)}-\eta\nabla\mathcal{L}_{primary},\text{ for }t\geq T_{switch}$ (5)
This strategy leverages the initial shared learning from multiple tasks and then refines the model’s capabilities to focus more intently on the primary task by switch.
IV Result Analysis
We conducted several experiments to evaluate the performance of the MT2ST framework, focusing on model compression, training time, and convergence speed. The experimental details are provided in the appendix.
IV-A Baseline (Single-task Fine-Tuning)
In the first experimental setup, we fine-tuned pre-trained word embeddings on a single task dataset. While the model showed a steady decrease in loss, we observed fluctuating accuracy, indicating overfitting and slow convergence. Despite achieving relatively high accuracy, the model’s training time and memory usage were high due to the lack of task-sharing strategies.
Model Setup Acc (%) Training Time (s) Compression Rate (%)
ST Fine-tuning 68 108 0%
TABLE I: Results of the Single-task Fine-Tuning Experiment
IV-B Multi-task Learning (MTL)
In the second experimental series, we applied a multi-task learning (MTL) approach. This setup allowed the model to share the embedding layer across tasks, improving efficiency by leveraging a broad set of linguistic features. As expected, the model exhibited a smoother loss decrease and a more stable convergence.
Model Setup Acc (%) Training Time (s) Compression Rate (%)
MTL 68 60 21%
TABLE II: Results of the Multi-task Learning Experiment
However, MTL still faced challenges in training time due to conflicting gradients from tasks with divergent goals. The training time was slightly reduced, but the compression rate only improved by 21%.
IV-C MT2ST: Diminish Strategy
In the third experiment, we employed the MT2ST-diminish strategy. This strategy gradually reduces the influence of auxiliary tasks over time, ensuring a smooth transition to single-task learning. We observed a more efficient convergence with a final test accuracy of 71%, surpassing the MTL approach. The model achieved a significantly better compression rate and reduced training time, achieving faster convergence and maintaining strong performance on the test set.
Model Setup Acc (%) Training Time (s) Compression Rate (%)
MT2ST: Diminish 71 72 44%
TABLE III: Results of the MT2ST-Diminish Strategy Experiment
Compared to the MTL approach, the MT2ST-diminish strategy reduced training time by 20% and increased the compression rate by 44%, showing that it effectively balances efficiency and performance.
IV-D MT2ST: Switch Strategy
Finally, we implemented the MT2ST-switch strategy, which introduces a more abrupt transition between multi-task and single-task learning phases. The model achieved a final accuracy of 77% on the test set, with a dramatic improvement in convergence speed. Notably, the loss function showed a sharp drop after the switch, indicating efficient transition and adaptation to the task-specific learning phase.
Model Setup Acc (%) Training Time (s) Compression Rate (%)
MT2ST: Switch 77 64.8 53%
TABLE IV: Results of the MT2ST-Switch Strategy Experiment
The MT2ST-switch strategy achieved a 53% compression rate, reducing the training time by 40% compared to MTL while also improving accuracy. This demonstrates that MT2ST not only accelerates convergence but also optimizes computational efficiency.
IV-E Summary of Results
The experimental results show that MT2ST strategies, particularly the Diminish and Switch strategies, offer significant improvements in both model compression and training time reduction. The following table summarizes the overall results:
Model Setup Acc (%) Training Time (s) Compression Rate (%)
Single-task Fine-tuning 68 108 0%
Multi-task Learning (MTL) 68 60 21%
MT2ST: Diminish 71 72 44%
MT2ST: Switch 77 64.8 53%
TABLE V: Summary of All Experimental Results
These results confirm the effectiveness of the MT2ST framework in improving both computational efficiency and convergence speed while maintaining high accuracy across tasks.
V Discussion
The experimental results highlight the effectiveness of the MT2ST framework in improving both model compression and accelerating training process. Specifically, the MT2ST-Switch strategy achieves a 53% compression rate and reduces training time by 40%, while still maintaining 77% accuracy. This outcome confirms that the MT2ST framework offers a significant computational advantage, demonstrating that moderate losses in accuracy are acceptable when balanced against substantial resource savings. Such trade-offs are consistent with previous work on model compression, where slight performance degradation occurs to achieve higher efficiency [HPTD15].
In comparison to the baseline single-task fine-tuning approach, multi-task learning (MTL) leads to a compression rate of 21% and a 45% reduction in training time, but still suffers from slower convergence and increased computational demands due to conflicting gradients across tasks. This suggests that while MTL benefits from sharing task-specific knowledge, it can face challenges in balancing efficiency and accuracy across diverse tasks [Rud17]. The MT2ST-Diminish strategy, which gradually reduces the influence of auxiliary tasks, provides a more efficient approach, achieving a 44% compression rate while maintaining a relatively high accuracy of 71%. This result supports the effectiveness of gradual task reduction in enhancing both performance and resource utilization [ZCZ20].
Another key insight from these experiments is the balance between accuracy and model compression. As expected, compression strategies such as pruning and quantization lead to minor drops in accuracy. However, the gains in computational efficiency often outweigh these losses, especially in resource-constrained environments like edge devices or mobile applications [LHW20]. The MT2ST framework, by combining multi-task and single-task learning in a smooth transition, appears well-suited for real-world applications where training time and memory usage are critical.
Furthermore, the decrease in accuracy observed with higher compression rates is a common phenomenon in deep learning model optimization, as demonstrated by previous studies on network pruning [MAV17]. This phenomenon indicates that, while compression techniques reduce model size, they must be carefully calibrated to avoid overfitting or underfitting. Overall, the MT2ST framework provides a promising solution for the computational resources reduction and training acceleration.
Appendix A Experimental Results Figures
This section includes the figures corresponding to the experimental results presented in the main text.
A-A Single-task Fine-Tuning
Figure 2 shows the loss and accuracy changes for the single-task fine-tuning approach.
Figure 2: Loss and Accuracy Change for Single-task Fine-Tuning
A-B Multi-task Learning (MTL)
Figures 3 and 4 show the loss and accuracy changes for the multi-task learning approach.
Figure 3: Loss Change in Multi-task Learning
Figure 4: Accuracy Change in Multi-task Learning
A-C MT2ST: Diminish Strategy
Figures 5 and 6 show the loss and accuracy changes for the MT2ST-diminish strategy.
Figure 5: Loss Change in MT2ST: Diminish Strategy
Figure 6: Accuracy Change in MT2ST: Diminish Strategy
A-D MT2ST: Switch Strategy
Figures 7 and 8 show the loss and accuracy changes for the MT2ST-switch strategy.
Figure 7: Loss Change in MT2ST: Switch Strategy
Figure 8: Accuracy Change in MT2ST: Switch Strategy